OpenStudy (albert0898):

Points P, N, and Z are collinear, AND NZ = 8, PN = 14, and PZ = 6. If point M is the midpoint of PN and point W is the midpoint of NZ, what is the length of MW? Explain it and please draw it out if possible.

OpenStudy (anonymous):

ok, because M is the midpoint of PN and PN=14 then PM=MN=7, same thing for the midpoint W, NZ=8 so WZ=NW=4 Now we want the length of MW MW= WZ-MZ WZ=4 MZ=1 then MW= 4-1=3